Circular No. 7439

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS D. A. Biesecker, Emergent Information Technologies, Inc., and God-

dard Space Flight Center, reports observations of thirteen additional Kreutz sungrazing comets detected by SOHO with the C2 telescope, with measurements by D. Hammer and Biesecker (the reductions and orbital elements by B. G. Marsden appearing on the MPECs specified below). None of the objects showed a tail. All of the comets were found via the SOHO website by M. Oates except for C/2000 L2, which was found by M. Boschat; Boschat and X. Leprette also found C/2000 L3.

1999 UT  α_{2000}  δ_{2000}  MPEC  Comet
May 5.326 2^{h}52.^{m}3 +14^{o}53′ 2000-L46  C/1999 J7
7.726 3 01.0 +15 29  2000-L47  C/1999 J8
9.226 3 06.6 +15 54  2000-L48  C/1999 J9
10.976 3 12.9 +16 17  2000-L49  C/1999 J10
14.504 3 26.6 +17 09  2000-L50  C/1999 J11
19.960 3 46.9 +18 29  2000-L51  C/1999 K11
23.563 3 59.9 +18 54  2000-L52  C/1999 K12
24.327 4 01.4 +18 53  2000-L53  C/1999 K13
27.976 4 17.6 +19 41  2000-L54  C/1999 K14
28.810 4 21.3 +19 49  2000-L55  C/1999 K15
2000 UT  α_{2000}  δ_{2000}  MPEC  Comet
June 3.479 4^{h}46.^{m}7 +20^{o}41′ 2000-L56  C/2000 L1
9.813 5 10.9 +21 10  2000-L57  C/2000 L2
10.165 5 12.8 +21 17  2000-L58  C/2000 L3

COMET C/1999 S4 (LINEAR) Visual m_{1} estimates: June 1.10 UT, 10.5 (C. Segarra, Algimia, Spain,

0.25-m reflector); 4.31, 10.4 (P. M. Raymundo, Salvador, Brazil, 0.25-m reflector); 6.75, 9.4 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector); 9.07, 9.4 (W. Hasubick, Buchloe, Germany, 25×100 binoculars); 11.98, 9.3 (A. Baransky, Kiev, Ukraine, 0.2-m reflector); 14.15, 8.9 (A. Pereira, Cabo da Roca, Portugal, 14×100 binoculars).

2000 June 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green